Graph the following: 2x+3y=6
Show step-by-step, please.

Answer by nyc_function(2741)

Your equation is linear. This means it
forms a straight line. All you need is
two points to graph a straight line.
Let x = 0 and solve for y.
2(0) + 3y = 6
0 + 3y = 6
3y = 6
y = 6/3
y = 2
The first point is (0,2).
Let y = 0 and solve for x
2x + 3(0) = 6
2x + 0 = 6
2x = 6
x = 6/2
x = 3
The second point to graph is (3,0).
Plot those two points on the coordinate plane
and connect them with a straight line.
Did you follow?